An impatient motorist considers speeding as he travels between two cities. If the trip normally takes 2.8 h at an average speed of 100.0 km/h, how much time will be saved (in minutes) if he exceeds the speed limit by 10.0 km/h?


1
Expert's answer
2021-02-22T10:20:10-0500

 If the trip takes 2.8 h at an average speed of 100.0 km/h, then the distance between these cities is:


"d = 100\dfrac{km}{h}\cdot 2.8h = 280km"

If the speed is "v = 100km\/h + 10km\/h = 110km\/h", then the time required to cover this distance is:


"t = \dfrac{d}{v}\\
t = \dfrac{280km}{110km/h} \approx 2.5h"

Answer. 2.5 h.